OpenStudy (anonymous):

Find all point of intersection of r=sin2q and r=sinq

OpenStudy (ash2326):

@nijokm We have \[r= \sin 2q\] and \[r =\sin q\] Equate the two to find the intersection points \[\sin 2q=\sin q\] \[\sin 2q-\sin q=0\] Now what's sin 2q in terms of sin q?

OpenStudy (ash2326):

@njiokm !!